OpenStudy (anonymous):

Use the given function value, and trigonometric identities (including the cofunction identities), to find the indicated trigonometric functions. sec θ = 5 (a) cos θ (b) cot θ (c) cot(90° − θ) (d) sin θ

OpenStudy (anonymous):

(a) use cos θ = 1 / sec θ ( b ) use cot θ = 1 / tan θ and sec^2 θ = 1 + tan^2 θ (d) use value of cos θ found in (a) in sin^2 θ = 1 - cos^2 θ

OpenStudy (anonymous):

oh yes (c) cot(90 - θ) = 1 / cot θ

OpenStudy (anonymous):

you ok with this?

OpenStudy (anonymous):

i'm told that part c is wrong

OpenStudy (anonymous):

hmm. - i did have to rack my brains with this one - long time... i'll check

OpenStudy (anonymous):

look at the diagram